OpenStudy (anonymous):

If y=2x^2-8x, what is the minimum value of the product xy?

OpenStudy (anonymous):

This is another problem from my practice AB exam I was given the other day that I don't quite remember how to do

OpenStudy (caominhim):

set the derivative equal to 0 then find the second derivative to see if it's a min or a max

OpenStudy (anonymous):

I thought the second derivative only tells you the concavity?

OpenStudy (anonymous):

The second derivative is just y=4

OpenStudy (caominhim):

you're right, scratch that second part

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The answer said that it was -8, but I am almost 100% sure that it is -16

OpenStudy (anonymous):

Because I set the first derivative= to 0 and I got x=2

OpenStudy (caominhim):

1 sec i'm checking with my roommate

OpenStudy (anonymous):

Then I plugged that into the equation for y and I got -8

OpenStudy (anonymous):

Ok thanks

OpenStudy (anonymous):

Then I multiplied -2 by 8 to get the product of xy

OpenStudy (caominhim):

okay so set derivative to 0....that's a critical point if at that point the 2nd derivative is positive that means that the first derivative just switched from negative to positive, signaling a min to find the y value plug the x value into the original equation

OpenStudy (anonymous):

Yes. I got 8 for the y val

OpenStudy (anonymous):

-8

OpenStudy (anonymous):

So I multiplied 2 by -8 and got -16. My stupid packet is telling me that the answer is -8

OpenStudy (anonymous):

I think that they forgot to multiply at the end tho

OpenStudy (caominhim):

i'm pretty sure it's -8, why are you sure it needs to be -16? 2(2)^2 -8(2) = -8

OpenStudy (anonymous):

That just gives you the y val

OpenStudy (anonymous):

They want the minimum product xy

OpenStudy (anonymous):

The minimum val of the product xy

OpenStudy (caominhim):

i never encountered a problem like that in calc, i think you may have interpreted the question wrong

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

The question is the beginning of the thread

OpenStudy (caominhim):

yea it asks for the product y, it doesn't say multiply x and y together

OpenStudy (anonymous):

I meant xy

OpenStudy (anonymous):

Sorry, that was a typo

OpenStudy (caominhim):

that doesn't make sense for this question, i took AB calc last year and never encountered a problem like this so I don't know how to help, sorry

OpenStudy (anonymous):

Its ok. Thanks anyways

OpenStudy (irishboy123):

"If y=2x^2-8x, what is the minimum value of the product xy?" P = xy = x (2x^2-8x) = ?? and then differeniate. easier?